# Simulations of Interacting Binary Systems - Pathways to Radio Bright GRB Progenitors

### v1

@@ 3.1.1 Dependence on Metallicity
The values of the final specific angular momentum of the massive star in cgs units range from to . This has important implications for the spin of the BH remnant when this massive star collapses, as well as the angular momentum in the disk that forms around the BH during/just after collapse. We consider this in our Discussion Section below.
3.1.2 Dependence on Mass Ratio
We examine the spin up of the massive star and change in orbital separation over the massive star lifetime for two mass ratios , shown in Fig. 2. For these simulations, we consider orbital periods of 1.5 and 3 days for a 25 massive star with a 10 and 15 BH, at stellar metallicities and . We find that the more extreme mass ratio encounters (with a 10 BH) lead to a larger spin up. This effect is more pronounced with smaller orbital periods, where stellar mass loss is greater due to the increase in encounter strength. This is consistent with results shown in de Mink et al. (2013) relating mass and angular momentum loss to mass ratio in interacting binaries.
However, we emphasize that for all of these encounters, the massive star is tidally distorted and there is a decrease in orbital separation of the system, leading to an increase in its spin angular momentum.
3.1.3 Dependence on Initial Spin of the Massive Star
We consider binary encounters where the massive star is initialized over a range of rotational velocities that include non-rotating , slowly rotating , and highly rotating stars. We use km/s from COSMIC output, as discussed in Kenoly et al. (2023). The break-up velocity is given by km/s for , respectively. Fig. 3 shows the stellar spin up and orbital separation at these varying initial rotation velocities, over the massive star lifetime. The evolution of and are similar for both cases of stellar metallicity. The spin for all three initial rotations reach close to the same value at the end of stellar lifetime. However, we note that when the massive star is initially rapidly rotating , this is the result of the massive star spinning down over its lifetime while increasing in orbital separation.
3.2 Stellar Density Profiles
In Figure 4, we show the density profile of the massive star at low and high metallicity ( and , respectively) for the case of an initial orbital period of three days. We expect that when tidal effects are not strong, earlier in the star’s life when the binary separation is still relatively large, the low metallicity star is more compact and the stellar density profiles are similar to single massive stars (black lines in Figure 4).
However, when the orbital separation decreases and tidal interactions become important, the companion can have an interesting effect on the density profile of the massive star (which in turn has important implications for the density profile of the circumbinary medium). As we see here, tidal interactions deform the star in such a way that it loses some mass and increases in radius with the likelihood, as discussed in Cheng & Evans (2013), for shock heating in the outer layers.
We expect that tidal interactions with an extended gas around a star may lead to ejection and therefore contribute to the density of the circumbinary environment. Provided that some of this gas settles into the direction into which the GRB jet (that is produced when the massive star collapses) propagates, this can lead to brighter radio emission since the radio flux scales as the square root of gas density333This is in the case of optically thin synchrotron emission. (Sari et al., 1998).
@@ 3.3 Additional Tests and Caveats
There are various mechanisms, such as the Spruit-Taylor dynamo, other magnetic coupling effects, or convection effects (Spruit, 2002; Heger et al., 2005), that can transport angular momentum out of the star throughout its lifetime. For example, Figure 3 of Heger et al. (2005) shows a pronounced decrease in angular momentum when magnetic fields are present, particularly for the pre-supernova stages of a star’s evolution. However, Denissenkov & Pinsonneault (2007) show that some of the assumptions required to invoke these effects are not always valid and, as a result, the loss of angular momentum can be significantly reduced. Potter et al. (2012) analyze the importance of the Spruit-Tayler dynamo in stars over a range of masses, and find that for stars above about 15 it becomes more difficult to sustain the dynamo, and as such the magnetic braking that normally occurs when this dynamo is active may less severe for more massive stars (particularly in the earlier stages of their evolution). We found that turning this flag on and off did not affect our results. We also investigated the effects of a rotating convection zone within the massive star. We began this investigation by setting the diffusion coefficients as prescribed in Heger et al. (2000) equal to 0. We find no change in the final angular momentum of the massive star at the end of its lifetime when this flag was turned on.
@@ 4 Discussion
We have investigated the properties of a massive star with a closely orbiting compact object companion and find a range of parameter space in which the massive star is spun up (i.e. ends its life with high angular momentum). We conjecture that this then is a plausible model for a highly spinning BH-disk system upon the massive star’s collapse, one that is capable of producing a viable GRB. We note that tidal synchronization is crucial for the massive star to spin up, especially if it begins with low initial angular momentum.
@@
for stellar mass , stellar radius , orbital separation , and where is the total mass of the binary system. The spin up of the massive star we find in our simulations in this work is consistent with this analytic result.
@@
We note that there is additional loss of orbital angular momentum through gravitational radiation (Peters, 1964) for which we have not accounted. However, for the orbital separations we consider (Newtonian gravity regime), this is a negligible effect and does not significantly affect our results over the timescales we are considering (Mapelli, 2020; Beradze et al., 2020). Other effects (that we do not consider in this study) such as common envelope evolution and dynamical friction can also change the angular momentum loss timescale, significantly affecting the orbital evolution and, in turn, the final angular momentum of the massive star.
@@
In connecting our progenitor system to a GRB - which results when the highly spinning massive star in our system collapses to a black hole - we have also not accounted for the potential loss of angular momentum upon collapse of the star. Here, we are making the assumption that a highly spinning massive star at the end of its life will produce a highly spinning BH-disk system. This is reasonably justified as shown in Jermyn et al. (2021a), who demonstrate that highly spinning massive stars in the ranges we consider end up with a roughly BH remnant, with a spin parameter , and with a long-lived disk (all elements to necessary to launch a relativistic GRB jet). Qin et al. (2018) use MESA to explore a range of conditions, in systems similar to the ones we consider here, where the spin of the second BH is high. They find that for very closely orbiting binaries, the spin of the second BH (i.e. the BH produced from the massive star in the systems we consider) is indeed near maximum (), particularly for low metallicity systems where mass loss from line driven winds is relatively small.
@@
We suggest that in some cases (especially for higher metallicity stars with a “puffier” density profile or for more vigorous tidal interactions), a denser (relative to a single massive star) medium may be formed around the binary system to a larger radius and may potentially lead to a brighter radio afterglow, where the radio flux scales as roughly the square root of the circumbinary density (Sari et al., 1998).
The tidal interaction between the black hole and the massive star is the underlying mechanism of the spin up of our massive star, as well as the extended density profile that deviates from the typical profile of a single massive star. We note that Sciarini et al. (2024) have pointed out that tidal prescriptions in binary stellar evolution codes can sometimes either over or underestimate the strength of the tidal torque and so caution must be taken in over-interpreting results.
@@ 5 Conclusions
We have modeled the evolution of a massive star with a closely orbiting compact object companion. Our study is motivated by the interesting result that radio bright GRBs tend to have a longer prompt gamma-ray burst duration compared to radio dark GRBs. As such, we have explored systems that potentially lead to a massive star with more angular momentum (and therefore a longer lived accretion disk and jet which reflects the duration of the prompt emission). We have also discussed the dependence of the density profile of the massive star on binary interaction and stellar metallicity (which has implications for the density of the circumbinary medium).
@@
We find that massive stars with closely orbiting compact object companions are significantly spun up through tidal interactions, and end their life with high angular momentum, conducive to producing a highly spinning BH-disk system that could power a relatively long-lived GRB jet. The specific angular momentum of the star at the end of its life aligns with the expectation of the angular momentum needed for a roughly 10 star to launch a relativistic jet through the Blanford-Znajek process (Blandford & Znajek, 1977).
We find tidal effects leading to significant stellar spin up for orbital periods up to about 4 days, and over a range of mass ratios.
We find that the tidal spin up is more pronounced for higher metallicity stars for a given binary orbital period, due to the larger radius/extent of the star. However, the spin up is still significantly present in low metallicity systems.
We find that spin up depends strongly on the initial rotation of the massive star. Unsurprisingly, stars with more initial angular momentum are not spun up for the orbital periods we consider. In these cases, tidal syncing causes the massive star to spin down to synchronize with the orbital period of the system.
We present two sample density profiles of a massive star at two metallicities, illustrating the greater extent of the higher metallicity star. We discuss how this not only affects tidal interactions but also potentially impacts the density profile of the medium surrounding the binary.
@@ Acknowledgments
We thank the MESA community for helpful conversations and the many resources provided to help run MESA. This work was supported by the U. S. Department of Energy through Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented was supported by the Laboratory Directed Research and Development program of LANL project number 20230115ER. We acknowledge LANL Institutional Computing HPC Resources under project w23extremex. Additional research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20210808PRD1. LA-UR-24-22983

### v2

@@ 3.3 Angular momentum
Notably, we identify the limit in parameter space which gives rise to a final increase in spin angular momentum for the parameter space of low metallicity () stars interacting with a BH on an initial twenty-one day orbital period. In comparing the second and third row of the first column of Fig. 3, we see an increase of without diminishment corresponding to mass loss due to RLOF that is less than mass loss due ot stellar wind. For initial orbital periods greater than twenty-one days, we find similar behavior with the magnitude of the peak decreasing with increasing initial orbital separation. The values of the final spin angular momentum of the massive star for simulations in Fig. 1 are given in CGS units. The corresponding values in final specific angular momentum range from to . This has important implications for the spin of the BH remnant when this massive star collapses, as well as the angular momentum in the disk that forms around the BH during/just after collapse. We consider this in our Discussion Section below.
Additionally, in the third row of Fig. 2 and Fig. 3, we show the change in spin angular momentum and the corresponding change in orbital angular momentum of the massive star in units of for each metallicity. In comparing these quantities, we find that for massive star – BH interactions with negligible mass loss, the decrease of matches the increase in . For binary interactions with significant mass loss and especially due to RLOF, we find coincident losses in the orbital angular momentum that deviate from the loss associated with the spin up of the massive star.
3.4 Termination conditions
We now discuss how we determine the final spin up and total mass loss of the massive star in our simulations of the interactions with a BH. Our conditions for termination, which set the final time for extracting these quantities, are based on tracking the formation of the helium and carbon-oxygen core as well as the central mass fraction of helium-4. For the sets of binary simulations corresponding to an initial three day (Fig. 2) and twenty-one day (Fig. 3) orbital period, we show in the fourth row the evolution of the formation of the helium () and carbon-oxygen () core in units of solar mass as well as the central mass fraction of helium-4 in units of stellar mass . We determine the final time of stellar termination as the coincident time where the helium an carbon-oxygen cores form and the central mass fraction of helium-4 decreases to . Furthermore, in the Appendix we discuss the correspondence between the formation of the carbon-oxygen core with an increase in central temperature. Using these conditions, present the final value of the change in spin angular momentum as well as the total mass loss in our simulations in Fig. 1 for binary parameter space listed in Tab. 1.
3.5 Dependence on Metallicity
We now discuss our results in terms of stellar metallicity in the low () and high () range of parameter space. In Fig. 1, we find that within initial orbital separations (periods of 20-50 days) where significant tidal effects increase the spin angular momentum, only stars with low metallicity contribute to the parameter space of viable GRB progenitors. This is primarily the case because of two reasons. First, the stellar radius is larger for massive stars with larger metallicities and therefore they are subject to stronger tidal interactions for a given orbital period than massive stars with smaller metallicities. Furthermore, when the stellar radius increases at the late stages of evolution, higher metallicity massive stars will be subject to greater RLOF. Second, the mass loss associated with stellar winds increases with stellar metallicity. This combined effect leads to significant mass loss for high metallicity stars and diminished spin angular momentum below the lower limit for viable GRB progenitors. In Fig. 2 and Fig. 3, we compare results between low and high metallicity stars for the initial three day and twenty-one day orbital periods. In the first row, we find the stellar to RLOF radius, , is closer to unity for high metallicity massive stars throughout its lifetime. In the second row, we see a corresponding greater mass loss due to both stellar wind and RLOF. In the third row, we find that this leads to greater diminishment in spin angular momentum.
3.6 Dependence on Mass Ratio
We consider binary interactions between a massive star and a and BH at mass ratios, , of 1.667 and 2.5, respectively. We find that the range in parameter space of viable GRB progenitors is similar at both mass ratios. In Fig. 1, we compare the final stellar spin up and total mass loss in binary systems with a (circle) and ("x") BH. We note that the spin-up of the star is smaller by a factor of for the smaller BH mass. The range of viable GRB candidates for low metallicity stars is within initial orbital periods of 20-40 days at both mass ratios.
3.7 Dependence on Initial Stellar Rotation
We consider the impact of initial rotation of the massive star in a binary with a 15 BH at a fixed initial orbital period. We compare simulation results initialized at three stellar velocities (, , ) under the critical limit of stellar break-up of a massive star at low () and high () metallicities. In Fig. 4 and Fig. 5, we show the massive star evolution at an initial orbital period of three and twenty-one days, respectively. Stellar age is given in units of years. In the first row, we compare the total mass loss . In the second row, we compare the ratio of stellar rotational velocity to initial orbital velocity . In the third row, we compare the change in stellar spin angular momentum with the change in orbital angular momentum in units of . In the fourth row, we track the evolution of the helium core mass in units of solar mass . For the initial orbital period of three days, in Fig. 4, we find similar total mass loss for all three initial rotations. For the initial orbital period of twenty-one days, in Fig. 5, there is greater total mass loss for the initial stellar rotation of in the low metallicity case, although at most on the order of of the initial mass of the star . Otherwise, the total mass loss is similar for all three initial rotations. For both cases in initial orbital period, we find that the stellar rotation of the massive star matches the initial rotation of the orbit at the end point of our simulations for all three initial rotations. In the second row of Fig. 4 and Fig. 5, we track the evolution of the ratio to and find that it approaches unity. Finally, in the third row of Fig. 4 and Fig. 5, we find the corresponding stellar spin angular momentum for all three initial rotations converges to a similar final spin-up value. This implies that in the binary parameters for stellar spin-up there is negligible dependence on initial stellar rotation because of tidal synchronization to the orbit in all cases for these initial orbital periods.
We consider the evolution of the massive star in radial profiles of stellar density and rotation. In Fig. 6, we compare profiles at early, middle, and late-stages of stellar evolution for an initially non-rotating 25 star at low metallicity () and high metallicity () at an initial orbital period of twenty-one days with a BH. In the top row, stellar density is given in CGS units versus stellar radius, given in solar radius . The bottom row shows the corresponding profile of stellar rotation in radians per second. At early stages, given by the solid line at stellar ages yrs, the stellar rotation profiles are close to numerical round-off error because the star is initially non-rotating. At middle stages, given by the dashed line at at yrs, we find that the massive star has expanded in radius close to a factor of two in comparison to early stages. This is consistent with the evolution of stellar radius in Fig. 3. Additionally, we note that for these times in Fig. 3 and Fig. 5, signficant mass loss through RLOF has yet to occur. The stellar rotation for both low and high metallicity is higher close to the center of the star than the outer layers. At late stages, given by the dashed dotted line at at yrs, we find that the massive star has signficantly expanded in radius close to a factor of ten in comparison to early stages. In comparing this time with Fig. 3 and Fig. 5, we note that this stellar age is characterized by RLOF and significant mass loss in the high metallicity case only. The stellar rotation for the low metallicity case is uniform across the radial profile of the star. The stellar rotation for high metallicity case shows higher rotation at the remnant core than the outer layers.
@@ 3.8 Additional Tests and Caveats
There are various mechanisms, such as the Spruit-Tayler dynamo, other magnetic coupling effects, or convection effects (Spruit, 2002; Heger et al., 2005; Moreno Méndez, 2014; Ma & Fuller, 2019), that can transport angular momentum out of the star throughout its lifetime. For example, Fig. 3 of Heger et al. (2005) shows a pronounced decrease in angular momentum when magnetic fields are present, particularly for the pre-supernova stages of a star’s evolution. However, Denissenkov & Pinsonneault (2007) show that some of the assumptions required to invoke these effects are not always valid and, as a result, the loss of angular momentum can be significantly reduced. Potter et al. (2012) analyze the importance of the Spruit-Tayler dynamo in stars over a range of masses, and find that for stars above about 15 it becomes more difficult to sustain the dynamo, and as such the magnetic braking that normally occurs when this dynamo is active may less severe for more massive stars (particularly in the earlier stages of their evolution). We learned that enabling or disabling this flag, while retaining the default MESA parameters, had no impact on our results. We also investigated the effects of a rotating convection zone within the massive star. We began this investigation by setting the diffusion coefficients as prescribed in Heger et al. (2000) to their default values according to MESA. We find no change in the final angular momentum of the massive star at the end of its lifetime when this flag was turned on.
@@ 4 Discussion
We have investigated the properties of a massive star with a closely orbiting BH companion and find a range of parameter space in which the massive star is spun up (i.e. ends its life with high angular momentum). We conjecture that this then is a plausible model for a highly spinning BH-disk system upon the massive star’s collapse, one that is capable of producing a viable GRB.
@@
for stellar mass , stellar radius , orbital separation , and where is the total mass of the binary system. The spin up of the massive star we find in our simulations in this work is consistent with this analytic result. We note that after collapse, the black hole central engine generally needs a spin parameter (ratio of its angular momentum to maximum possible angular momentum) of at least about 0.5 to launch a jet through the BZ process. For a five solar mass black hole this corresponds to roughly g cm. For our low metallicity simulations in particular, we find that the star at the end of its life has angular momentum in the range of to g cm.
Our full range of binary parameter space consideration may be compared to estimates from perturbation theory which characterize the tidal encounter through a strength parameter (Press & Teukolsky, 1977), for circular orbits at stellar-BH separation distance ,
for orbital timescale and stellar dynamical timescale . For a given star, encounters with smaller orbital periods will give rise to smaller or stronger tidal interactions. Note that the limit corresponds to the tidal disruption limit where the dynamical time scales for the orbit and star are comparable. In our study, we consider the weak tidal interaction regime and provide detailed simulations using the stellar evolution model described in Sec. 2, that probe the non-linear, hydrodynamic response of a massive star throughout its lifetime. We find that the increase spin up of the massive star due to a decrease in orbital period are consistent with results of the spin up of polytropic stars under short time-scale interactions in parabolic orbits (Kochanek, 1992; Khokhlov et al., 1993; Cheng & Evans, 2013).
However, we find that modeling the long-term interaction between a massive star and BH across stellar lifetime leads to the identification of a parameter region of diminished, final stellar spin angular momentum due to mass loss driven by RLOF as the stellar radius expands prior to collapse. These results which show a sharp decrease in stellar spin coincident with an increase in mass loss in late-stage stellar evolution. Paxton et al. (2013) discuss the challenges and uncertainty in modeling the radiation-pressure dominated envelopes in massive stars at late phases in the evolution with a 1D stellar evolution code. Nonetheless, we present our results at stellar termination with MESA while noting that follow-up numerical investigations that go beyond its limitations are necessary to resolve these uncertainties.
@@
Regarding the parameter space of encounters with competing effects that change the stellar angular momentum, Detmers et al. (2008) find that the tidal interaction leads to significant spin-up as well as significant losses due to mass loss due to winds and RLOF between Wolf-Rayet stars at solar metallicity and compact object companions with initial orbital periods less than 24 hours. We note that there is additional loss of orbital angular momentum through gravitational radiation (Peters, 1964) for which we have not accounted. However, for the orbital separations we consider (Newtonian gravity regime), this is a negligible effect and does not significantly affect our results over the timescales we are considering (Mapelli, 2020; Beradze et al., 2020). Other effects (that we do not consider in this study) such as common envelope evolution and dynamical friction can also change the angular momentum loss timescale, significantly affecting the orbital evolution and, in turn, the final angular momentum of the massive star.
@@
In connecting our progenitor system to a GRB - which results when the highly spinning massive star in our system collapses to a BH - we have also not accounted for the potential loss of angular momentum upon collapse of the star. Here, we are making the assumption that a highly spinning massive star at the end of its life will produce a highly spinning BH-disk system. This is reasonably justified as shown in Jermyn et al. (2021a), who demonstrate that highly spinning massive stars in the ranges we consider end up with a roughly BH remnant, with a spin parameter , and with a long-lived disk (all elements to necessary to launch a relativistic GRB jet). Qin et al. (2018) use MESA to explore a range of conditions, in systems similar to the ones we consider here, where the spin of the second BH is high. They find that for very closely orbiting binaries, the spin of the second BH (i.e. the BH produced from the massive star in the systems we consider) is indeed near maximum (), particularly for low metallicity systems where mass loss from line driven winds is relatively small. Ma & Fuller (2023) use MESA to investigate the interaction between a WR star with a BH companion. They find that a WR star at solar metallicity experiences significant angular momentum loss due to stellar winds. In contrast, WR stars with low metallicity retain their angular momentum and are spun up due to tidal interactions.
@@
We suggest that in some cases (especially for higher metallicity stars with a “puffier” density profile or for more vigorous tidal interactions), a denser (relative to a single massive star) medium may be formed around the binary system to a larger radius and may potentially lead to a brighter radio afterglow, where the radio flux scales as roughly the square root of the circumbinary density (Sari et al., 1998). The specifics of the mass-loss from line driven winds in these systems will also affect the circumbinary density profile. In this paper, we simply offer the suggestion of a varying circumbinary environment depending on the resulting density profiles of star, as seen in Figure 6, but leave a detailed investigation of this issue to a following publication.
The tidal interaction between the BH and the massive star is the underlying mechanism of the spin up of our massive star, as well as the extended density profile that deviates from the typical profile of a single massive star. We note that Sciarini et al. (2024) have pointed out that tidal prescriptions in binary stellar evolution codes can sometimes either over or underestimate the strength of the tidal torque and so caution must be taken in over-interpreting results.
@@ 5 Conclusions
In this paper we have explored the influence of a BH companion on a massive star throughout its lifetime, over a range of parameter space varying orbital period, metallicity, initial spin of the star, mass ratio, accretion rate, magnetic fields, and numerical stopping conditions. Our study is motivated by the interesting result that radio bright GRBs tend to have a longer prompt gamma-ray burst duration compared to radio dark GRBs. As such, we have explored systems that potentially lead to a massive star with more angular momentum (and therefore a longer lived accretion disk and jet which reflects the duration of the prompt emission). We have focused on a range of parameter space in which the initial orbital period (binary separation) is such that the massive star does not initially experience RLOF and tidal interactions still play a significant role in the spin-angular momentum evolution of the massive star. We have also briefly discussed the dependence of the density profile of the massive star on binary interaction and stellar metallicity (which has implications for the density of the circumbinary medium).
@@
We identify a particular range of parameter space where massive stars with closely orbiting BH companions are significantly spun up through tidal interactions, and end their life with high angular momentum, conducive to producing a highly spinning BH-disk system that could power a relatively long-lived GRB jet. The specific angular momentum of the star at the end of its life aligns with the expectation of the angular momentum needed for a roughly 10 star to launch a relativistic jet through the Blanford-Znajek process (Blandford & Znajek, 1977).
We identify low metallicity () massive stars interacting with both 15 and 10 BH for an initial orbital period of 20-50 days as viable progenitor binary systems. For this range of parameter space, we note significant final stellar spin-up and negligible final total mass loss on the order of a factor of of the initial mass of the star. Outside of this region, non-viable binaries are characterized by significant final mass loss close to a factor of unity or at an orbital separation where the tidal interaction is negligible ( 60). Additionally, we note a very slight decrease in orbital separation on the order of for simulations with negligible mass loss including the viable progenitor binary systems.
We find that the final stellar spin angular momentum does not depend on the initial rotation of the massive star for the parameters we have investigated. In these cases, the tidal effects will spin up or spin down the massive star in order to synchronize with the orbital period of the system.
@@ Acknowledgments
We thank the referee for the detailed and constructive feedback. We thank the MESA community for helpful conversations and the many resources provided to help run MESA. This work was supported by the U. S. Department of Energy through Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). Research presented was supported by the Laboratory Directed Research and Development program of LANL project number 20230115ER. We acknowledge LANL Institutional Computing HPC Resources under project w23extremex. Additional research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project number 20210808PRD1. LA-UR-24-22983